OpenStudy (anonymous):

One end of a 30-cm long aluminium rod is exposed to a temperature of 500oC while the other end is maintained at 20oC. The rod has the diameter of 2.5 cm. If heat is conducted through the rod at the rate of 164.9J/s, calculate the the thermal conductivity of aluminium.

OpenStudy (anonymous):

\[\large H=\frac{\Delta Q}{\Delta t} = k A\frac{\Delta T}{x}\] where \[\frac{\Delta Q}{\Delta t}\] -->Rate of flow of heat k is the thermal conductivity, A is the total cross sectional area of conducting surface, ΔT is temperature difference, and x is the thickness of conducting surface separating the two temperatures.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome :D